Give an example of a function of two variables whose limit at (0

,

0) does not exist, that is

lim

(

x,y

)

(0

,

0)

f

(

x, y

) does not exist. Explain also why the limit does not exist.



1
Expert's answer
2021-11-25T23:58:30-0500

Question: Give an example of a function of two variables whose limit at (0,0) does not exist, that is lim(x,y)→(0,0).

f(x, y) does not exist. Explain also why the limit does not exist.


Answer:

"lim_{\n(x,y)\u2192(0,0)}\n\u200b\n \n\\dfrac{x\u00b2\u2212y\u00b2}\n{x\u00b2+y\u00b2}\n\u200b"

The above function does not exist.

This is an example of a function of two variables whose limit at (0, 0)

The reason of non existence of the limit is;

If y = mx, and "y \to 0, x \to 0" for any arbitrary m. The limit becomes

"lim_{x\to0}\dfrac{x²-m²x²}{x²+m²x²}
 
\\
= lim_{x\to0}\dfrac{1-m²}{1+m²}\\

= \dfrac{1-m²}{1+m²}"


So the value of the limit varies as the value of m varies.

Hence the limit does not exist"."